OpenStudy (anonymous):

Write the equation of the line that passes through the points (5,2) and (-3,2)

OpenStudy (anonymous):

Equation of a line: y - y1 = m(x-x1) You need to find m, by calculating y2-y1/x2-x1, where y2 = 2, y1 = 2, x2 = -3, x1 = 5

OpenStudy (anonymous):

m is the gradient, by the way.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

how did you get the 2 ?

OpenStudy (anonymous):

oh okay thank you ! :)